OpenStudy (anonymous):

solve for x 132 = 1 over 2x + 2 (x - 9)

OpenStudy (anonymous):

\[132 =\frac{ 1 }{ 2x + 2(x-9) }\]

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

so how is the equation supposed to look like please clarify :)

OpenStudy (anonymous):

its 132 = 1/2x + 2 (x - 9 )

OpenStudy (anonymous):

132 =  x +2 ( x - 9 )

OpenStudy (jiteshmeghwal9):

\[\LARGE{132=\frac{1}{2x}+2(x-9)}\]

OpenStudy (anonymous):

Use parentheses when writing your equation. Otherwise there could be a number of different interpretations.\[132 = \frac{1}{2}x + 2(x-9)\] could be written 132=(1/2)x+2(x-9) Or is it \[132 = \frac{1}{2x} + 2(x-9)\]?

OpenStudy (anonymous):

ah ok now I understand better. So first distribute 2 to x - 9 2x - 18 132 = 1/2x + 2x - 18 Then combine like terms which is 1/2x + 2x 132 = 2 1/2x - 18. Then add 18 to both sides. +18 +18 --------------------- 150 = 2 1/2x Then you divide. I got 60! Hope this helped!

OpenStudy (anonymous):

thanks